I have a problem in solving the following question, Can some one provide me the detailed solution of it,


The moment of inertia of a uniform semicircular disc of mass M and radius r about a line perpendicular to the plane of the disc through the centre is ...?

Grade:12

4 Answers

dev bhatnagar
15 years ago

This a simple conceptual question I think following can help you

Mass of semicircular disc = M

Suppose there is a circular disc of mass 2M, then

Moment of intertia of circular disc = 1/2(2M)R2

Moment of intertia of circular disc = 1/2(2M)R2 = MR2

=> So, Moment of intertia of semi-circular disc = (1/2)MR2

sowmya thandra
15 years ago

can i get to know the detail information abt calculating moment of inertia

Tripti Arora
6 years ago
The above answer is right but from wrong method because moment of inertia is same for any arc or any part of disc it is not that if disc is half then mass is also half and Moi also half
parth
5 years ago
As it is uniform the mass didtribution is same all over the disc so half of the disc results in ½ of mass
 
  • m=density(d)*vol
  • M1=d*t*\displaystyle \piR2
  • M2=d*t*\displaystyle \pi(r/2)2